# Advancing Reliable Test-Time Adaptation of Vision-Language Models under Visual Variations

### v1

@@ Abstract.
Vision-language models (VLMs) exhibit remarkable zero-shot capabilities but struggle with distribution shifts in downstream tasks when labeled data is unavailable, which has motivated the development of Test-Time Adaptation (TTA) to improve VLMs’ performance during inference without annotations. Among various TTA approaches, cache-based methods show promise by preserving historical knowledge from low-entropy samples in a dynamic cache and fostering efficient adaptation. However, these methods face two critical reliability challenges: (1) entropy often becomes unreliable under distribution shifts, causing error accumulation in the cache and degradation in adaptation performance; (2) the final predictions may be unreliable due to inflexible decision boundaries that fail to accommodate large downstream shifts. To address these challenges, we propose a Reliable Test-time Adaptation (ReTA) method that integrates two complementary strategies to enhance reliability from two perspectives. First, to mitigate the unreliability of entropy as a sample selection criterion for cache construction, we introduce Consistency-aware Entropy Reweighting (CER), which incorporates consistency constraints to weight entropy during cache updating. While conventional approaches rely solely on low entropy for cache prioritization and risk introducing noise, our method leverages predictive consistency to maintain a high-quality cache and facilitate more robust adaptation. Second, we present Diversity-driven Distribution Calibration (DDC), which models class-wise text embeddings as multivariate Gaussian distributions, enabling adaptive decision boundaries for more accurate predictions across visually diverse content. Extensive experiments demonstrate that ReTA consistently outperforms state-of-the-art methods, particularly under challenging real-world distribution shifts.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06submissionid: 4740ccs: Computing methodologies Transfer learning
@@ 1. Introduction
Large-scale vision-language models (VLMs), such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021), pre-trained on massive web-scale datasets, have demonstrated impressive zero-shot capabilities and strong open-world visual understanding across a wide range of vision tasks, including classification (Sun et al., 2022; Ding et al., 2023, 2025), retrieval (Xie et al., 2023; Iscen et al., 2024), and segmentation (Jiahaoli et al., 2024; Zhu et al., 2023). Their core mechanism involves learning a shared representation space by aligning visual and textual modalities during training, and subsequently employing similarity-based matching for classification at inference. However, VLMs often suffer from performance degradation when deployed on unlabeled test data from downstream domains that exhibit significant distribution shifts from their pre-training distribution (Agarwal et al., 2021; Menon et al., 2024; Parashar et al., 2024), limiting their effectiveness in real-world scenarios with diverse visual variations.
@@
  • We introduce Consistency-aware Entropy Reweighting (CER), which constructs class-specific textual representations to evaluate prediction consistency across adjacent semantic representations. CER identifies high-consistency samples and prioritizes them during cache updates to improve adaptation quality.
@@
  • Our comprehensive evaluation across multiple benchmarks demonstrates that ReTA consistently outperforms state-of-the-art TTA methods, confirming its effectiveness in enhancing VLM adaptation.
@@ 3.1. Preliminaries
where $\mathbf{F}_{\text{cache }}$ is the stored image feature, $\mathbf{L}_{p}$ is a one-hot pseudo-label obtained from CLIP prediction, and the entropy $H(p)$ is defined as:
(3) $H(p)=\left(-\sum_{c=1}^{C}p^{c}_{\scriptsize\text{CLIP}}\log p^{c}_{\scriptsize\text{CLIP}}\right).$
@@ 3.2. Consistency-aware Entropy Reweighting
Previous methods (Zhang et al., 2024a; Karmanov et al., 2024; Zhang et al., 2024b) take low-entropy cached samples as visual prototypes for feature retrieval and cache prediction computation. However, relying solely on entropy is unreliable as CLIP is usually overconfident, potentially assigning high confidence to misclassified samples (Farina et al., [n. d.]; Khan and Fu, 2024a). Since prediction confidence lacks correlation with pseudo-label accuracy (Huang et al., 2022; Xiong et al., 2023), low-entropy samples may carry incorrect labels, resulting in cached prototypes deviating from the expected distribution and degrading retrieval performance.
While approaches like TDA (Karmanov et al., 2024) and DPE (Zhang et al., 2024a) have incorporated knowledge-enriched text embeddings with different prompts, they typically average them into a single representation, which fails to fully exploit their semantic diversity. To address this limitation, we propose Consistency-aware Entropy Reweighting (CER), which improves cache quality by prioritizing the storage of relatively reliable samples. Specifically, CER leverages prediction consistency across neighboring semantic representations to identify reliable and representative samples. Inspired by the committee-based strategies (Lyu et al., 2023; Wang et al., 2017; Seung et al., 1992), we construct a semantic voting committee using multiple text embeddings to assess prediction reliability. While controversial samples expose uncertainty in committee members’ judgments, CER prioritizes those with consistent agreement in voting, as they are more representative and reliable for caching.
@@
We then sort text embeddings by relative cosine similarity in ascending order, arranging them from semantic outliers to centroids, and construct our final adjacent text embedding set $\left\{\widehat{\boldsymbol{t}}_{1}^{c},\widehat{\boldsymbol{t}}_{2}^{c},\ldots,\widehat{\boldsymbol{t}}_{M}^{c}\right\}$ with $M$ embeddings per class using a progressive binning approach. Specifically, we define $Q_{m}=\lfloor\frac{mK}{M}\rfloor$ for $m=1,2,...,M$ to determine embeddings for the $m$-th adjacent embedding. Each $\widehat{\boldsymbol{t}}_{m}^{c}$ is created by averaging the first $Q_{m}$ embeddings from the sorted sequence, where embeddings are cumulatively included in progressively expanding pools. Our ascending progressive binning strategy creates a smooth semantic transition across embeddings while preserving diversity, thus forming a natural semantic neighborhood for subsequent consistency assessment. Functionally, these adjacent text embeddings operate as a semantic voting committee, where each embedding offers a diverse perspective on class prediction.
@@
where $y=\arg\max\nolimits_{c\in{C}}(\mathbf{z}^{T}\widehat{\boldsymbol{t}}_{M}^{c})$ denotes the original prediction based on the image feature and the final text prototype, and $\gamma>1$ is a penalty parameter that increases entropy, thereby reducing the priority of inconsistent predictions. Additionally, we define a stability factor $\mathcal{S}$:
[…]
@@
3.3. Diversity-driven Distribution Calibration
While CER provides reliable features for cache-based adaptation, it remains challenging to make rapid adjustments to each test-time input, as the cache contains only a few high-confidence samples and may be limited in accommodating visual variations. Existing cache-based methods (Karmanov et al., 2024; Zhang et al., 2024b, c) attempt to compensate for this drawback by enhancing discriminability through sophisticated prompts, but their adaptation capabilities remain limited in dynamic environments due to VLMs’ fixed decision boundaries, which struggle to handle distribution shifts. To overcome this challenge, we propose Diversity-driven Distribution Calibration (DDC), which leverages adjacent text embeddings from CER to construct a dynamically adjustable distribution. Drawing from recent success in modeling prompt embeddings with Gaussian distributions (Lu et al., 2022; Bao et al., 2025), we characterize the adjacent text embeddings $\tilde{\mathcal{T}}$ as an approximate Gaussian distribution $\mathcal{N}\left(\boldsymbol{\mu}^{c},\boldsymbol{\Sigma}^{c}\right)$, where $\boldsymbol{\mu}^{c}$ representing the mean embedding of class $c$ and $\boldsymbol{\Sigma}^{c}$ denoting the corresponding covariance matrix. Building on the efficient residual learning strategy in DPE (Zhang et al., 2024a), we extend this idea to adapt flexible representational distributions, enabling DDC to accommodate evolving visual features.
(16) $\widehat{\boldsymbol{t}}_{m}^{c}=\frac{\widehat{\boldsymbol{t}}_{m}^{c}+\boldsymbol{r}_{m}^{c}}{||\widehat{\boldsymbol{t}}_{m}^{c}+\boldsymbol{r}_{m}^{c}||}$
(17) $\widehat{\boldsymbol{t}}_{m}^{c}=\frac{(l-1)\widehat{\boldsymbol{t}}_{m}^{c}+\widehat{\boldsymbol{t}}_{m}^{c(\ast)}}{||(l-1)\widehat{\boldsymbol{t}}_{m}^{c}+\widehat{\boldsymbol{t}}_{m}^{c(\ast)}||}$
@@ 3.3. Diversity-driven Distribution Calibration
$&\mathcal{L}_{\text{ent}}=H(\tilde{p}_{\text{cls}}(\mathbf{x}))=-\sum_{c=1}^{C}\tilde{p}_{\text{cls}}^{c}\log(\tilde{p}_{\text{cls}}^{c})\\ &\text{with}\;\tilde{p}^{\text{cls}}_{c}=\frac{\sum_{i=1}^{N}\mathbb{I}[H(p_{\text{cls}}^{c}(\mathbf{x}_{i}))<\delta]\cdot p_{\text{cls}}^{c}(\mathbf{x}_{i})}{\sum_{i=1}^{N}\mathbb{I}[H(p_{\text{cls}}^{c}(\mathbf{x}_{i}))<\delta]}\end{aligned}$
@@
$\mathcal{L}_{\text{surr}}&=\underset{(\mathbf{z},\tilde{y})}{\mathbb{E}}\left[-\log\frac{\exp\left(\left\langle\mathbf{z},\widehat{\boldsymbol{t}}_{M}^{\tilde{y}}\right\rangle/\tau\right)}{\sum_{c=1}^{C}\exp\left(\left\langle\mathbf{z},\widehat{\boldsymbol{t}}_{M}^{c}\right\rangle/\tau+\mathbf{z}^{T}\mathbf{W}_{c,\tilde{y}}\mathbf{z}/2\tau^{2}\right)}\right]\end{aligned}$
@@ 3.4. Unified Learning Process and Inference
4. Experiments
4.1. Experimental Setup
4.2. Evaluation of Cross-Datasets Generalization
@@ 4.4. Experimental Analysis
Figure 4. Comparison of cache accuracy and overall performance across four datasets.
@@
Comparison of Cache Reliability and Final Performance. Figure 4 compares cache accuracy and overall classification performance across four fine-grained datasets. ReTA shows remarkable performance in both metrics, achieving 74.52% on UCF101 (surpassing TDA by 3.86% and DPE by 4.08%) and 77.55% on Flowers (exceeding TDA by 6.13% and DPE by 2.48%). Combined with evidence from Figure 1, these results prove that ReTA achieves superior final accuracy while maintaining high cache reliability throughout adaptation.
@@
Method Testing Time Acc. Gain

### v2

@@ Abstract.
Vision-language models (VLMs) exhibit remarkable zero-shot capabilities but struggle with distribution shifts in downstream tasks when labeled data is unavailable, which has motivated the development of Test-Time Adaptation (TTA) to improve VLMs’ performance during inference without annotations. Among various TTA approaches, cache-based methods show promise by preserving historical knowledge from low-entropy samples in a dynamic cache and fostering efficient adaptation. However, these methods face two critical reliability challenges: (1) entropy often becomes unreliable under distribution shifts, causing error accumulation in the cache and degradation in adaptation performance; (2) the final predictions may be unreliable due to inflexible decision boundaries that fail to accommodate large downstream shifts. To address these challenges, we propose a Reliable Test-time Adaptation (ReTA) [^1] Code is available at https://github.com/Evelyn1ywliang/ReTA. method that integrates two complementary strategies to enhance reliability from two perspectives. First, to mitigate the unreliability of entropy as a sample selection criterion for cache construction, we introduce Consistency-aware Entropy Reweighting (CER), which incorporates consistency constraints to weight entropy during cache updating. While conventional approaches rely solely on low entropy for cache prioritization and risk introducing noise, our method leverages predictive consistency to maintain a high-quality cache and facilitate more robust adaptation. Second, we present Diversity-driven Distribution Calibration (DDC), which models class-wise text embeddings as multivariate Gaussian distributions, enabling adaptive decision boundaries for more accurate predictions across visually diverse content. Extensive experiments demonstrate that ReTA consistently outperforms state-of-the-art methods, particularly under real-world distribution shifts.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755607isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Transfer learning
@@ 1. Introduction
Large-scale vision-language models (VLMs), such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021), pre-trained on massive web-scale datasets, have demonstrated impressive zero-shot capabilities and strong open-world visual understanding across a wide range of vision tasks, including classification (Ding et al., 2023; Wang et al., 2023a, 2025), retrieval (Xie et al., 2023; Iscen et al., 2024), and segmentation (Jiahaoli et al., 2024; Hao et al., 2023). Their core mechanism involves learning a shared representation space by aligning visual and textual modalities during training, and subsequently employing similarity-based matching for classification at inference. However, VLMs often suffer from performance degradation when deployed on unlabeled test data from downstream domains that exhibit significant distribution shifts from their pre-training distribution (Agarwal et al., 2021; Lyu et al., 2025; Xiong et al., 2025), limiting their effectiveness in real-world scenarios with diverse visual variations.
@@
  • We introduce Consistency-aware Entropy Reweighting (CER), which constructs class-specific textual representations to evaluate prediction consistency across adjacent representations. CER identifies high-consistency samples and prioritizes them during cache updates to enhance adaptation.
@@
  • Our comprehensive evaluation across multiple benchmarks demonstrates that ReTA consistently outperforms state-of-the-art TTA methods and is effective in adapting VLMs.
@@ 3.1. Preliminaries
[…]
$H(p)=\left(-\sum_{c=1}^{C}p^{c}_{\scriptsize\text{CLIP}}\log p^{c}_{\scriptsize\text{CLIP}}\right).$
@@ 3.2. Consistency-aware Entropy Reweighting
Previous methods (Zhang et al., 2024a; Karmanov et al., 2024; Zhang et al., 2024b) take low-entropy cached samples as visual prototypes for feature retrieval and cache prediction computation. However, relying solely on entropy is unreliable as CLIP is usually overconfident, potentially assigning high confidence to misclassified samples (Farina et al., [n. d.]; Khan and Fu, 2024a). Since prediction confidence lacks correlation with pseudo-label accuracy (Huang et al., 2022; Xiong et al., 2023), low-entropy samples may carry incorrect labels, resulting in cached prototypes deviating from the expected distribution and degrading retrieval performance. While approaches like TDA (Karmanov et al., 2024) and DPE (Zhang et al., 2024a) have incorporated knowledge-enriched text embeddings with different prompts, they typically average them into a single representation, which fails to fully exploit their semantic diversity.
To address this limitation, we propose Consistency-aware Entropy Reweighting (CER), which improves cache quality by prioritizing the storage of relatively reliable samples. Specifically, CER leverages prediction consistency across neighboring semantic representations to identify reliable and representative samples. Inspired by the committee-based strategies (Lyu et al., 2023; Wang et al., 2017; Seung et al., 1992), we construct a semantic voting committee using multiple text embeddings to assess prediction reliability. While controversial samples expose uncertainty in committee members’ judgments, CER prioritizes those with consistent agreement in voting, as they are more representative and reliable for caching.
@@
We then sort text embeddings by relative cosine similarity in ascending order, arranging them from semantic outliers to centroids, and construct our final adjacent text embedding set $\left\{\widehat{\boldsymbol{t}}_{1}^{c},\widehat{\boldsymbol{t}}_{2}^{c},\ldots,\widehat{\boldsymbol{t}}_{M}^{c}\right\}$ with $M$ embeddings per class using a progressive binning approach. Specifically, we define $Q_{m}=\lfloor\frac{mK}{M}\rfloor$ for m=1,2,,Mm=1,2,...,M to determine embeddings for the $m$-th adjacent embedding. Each $\widehat{\boldsymbol{t}}_{m}^{c}$ is created by averaging the first $Q_{m}$ embeddings from the sorted sequence, where embeddings are cumulatively included in progressively expanding pools. Our ascending progressive binning strategy creates a smooth semantic transition across embeddings while preserving diversity, thus forming a natural semantic neighborhood for subsequent consistency assessment. Functionally, these adjacent text embeddings operate as a semantic voting committee, where each embedding offers a diverse perspective on class prediction.
@@
where $y=\arg\max\nolimits_{c\in{C}}(\mathbf{z}^{T}\widehat{\boldsymbol{t}}_{M}^{c})$ denotes the original prediction based on $\mathbf{z}$ and the final text prototype, and $\gamma>1$ is a penalty parameter that increases entropy, thereby reducing the priority of inconsistent predictions. Additionally, we define a stability factor $\mathcal{S}$:
(12) $\text{Stability: }\mathcal{S}=\frac{M}{n^{\ast}},\quad 1\leq\mathcal{S}\leq M$
@@
3.3. Diversity-driven Distribution Calibration
While CER provides reliable features for cache-based adaptation, it remains challenging to make rapid adjustments to each test-time input, as the cache contains only a few high-confidence samples and may be limited in accommodating visual variations. Existing cache-based methods (Karmanov et al., 2024; Zhang et al., 2024b, c) attempt to compensate for this drawback by enhancing discriminability through sophisticated prompts, but their adaptation capabilities remain limited in dynamic environments due to VLMs’ fixed decision boundaries, which struggle to handle distribution shifts. To overcome this challenge, we propose Diversity-driven Distribution Calibration (DDC), which leverages adjacent text embeddings from CER to construct a dynamically adjustable distribution. Drawing from recent success in modeling prompt embeddings with Gaussian distributions (Lu et al., 2022; Bao et al., 2025a), we characterize the adjacent text embeddings $\tilde{\mathcal{T}}$ as an approximate Gaussian distribution $\mathcal{N}\left(\boldsymbol{\mu}^{c},\boldsymbol{\Sigma}^{c}\right)$, where $\boldsymbol{\mu}^{c}$ representing the mean embedding of class $c$ and $\boldsymbol{\Sigma}^{c}$ denoting the corresponding covariance matrix. Building on the efficient residual learning strategy in DPE (Zhang et al., 2024a), we extend this idea to adapt flexible representational distributions, enabling DDC to accommodate evolving visual features.
[…]
@@ 3.3. Diversity-driven Distribution Calibration
$\begin{aligned} &\mathcal{L}_{\text{ent}}=H(\tilde{p}_{\text{cls}}(\mathbf{x}))=-\sum_{c=1}^{C}\tilde{p}_{\text{cls}}^{c}\log(\tilde{p}_{\text{cls}}^{c})\\ &\text{with}\;\tilde{p}^{\text{cls}}_{c}=\frac{\sum_{i=1}^{N}\mathbb{I}[H(p_{\text{cls}}^{c}(\mathbf{x}_{i}))<\delta]\cdot p_{\text{cls}}^{c}(\mathbf{x}_{i})}{\sum_{i=1}^{N}\mathbb{I}[H(p_{\text{cls}}^{c}(\mathbf{x}_{i}))<\delta]}\end{aligned}$
@@
$\begin{aligned} \mathcal{L}_{\text{surr}}&=\underset{(\mathbf{z},\tilde{y})}{\mathbb{E}}\left[-\log\frac{\exp\left(\left\langle\mathbf{z},\widehat{\boldsymbol{t}}_{M}^{\tilde{y}}\right\rangle/\tau\right)}{\sum_{c=1}^{C}\exp\left(\left\langle\mathbf{z},\widehat{\boldsymbol{t}}_{M}^{c}\right\rangle/\tau+\mathbf{z}^{T}\mathbf{W}_{c,\tilde{y}}\mathbf{z}/2\tau^{2}\right)}\right]\end{aligned}$
@@ 3.4. Unified Learning Process and Inference
4. Experiments
4.1. Experimental Setup
4.2. Evaluation of Cross-Datasets Generalization
@@ 4.4. Experimental Analysis
Figure 4. Final comparison of adaptation performance and cache accuracy across four datasets for cache-based methods.
@@
Comparison of Cache Reliability and Final Performance. Figure 4 compares cache accuracy and overall classification performance across four fine-grained datasets. ReTA shows remarkable performance in both metrics, achieving 74.52% on UCF101 (surpassing TDA by 3.86% and DPE by 4.08%) and 77.55% on Flowers (exceeding TDA by 6.13% and DPE by 2.48%). Combined with evidence from Figure 1, these results prove that ReTA achieves superior final accuracy while maintaining high cache reliability.
@@
Method
Testing Time
Acc. Gain